OpenStudy (anonymous):

6x^2-2x-1=0 how can i get this using a quadratic equation

OpenStudy (anonymous):

Do you mean Quadratic formula? If so, let a = 6 b = -2 what is c?

OpenStudy (anonymous):

a = 6 b = -2 c = -1 \[-b \pm \sqrt{b^2-4ac} \over 2a\]

OpenStudy (anonymous):

Apply a quadratic of the form: \[\Large a{x^2} + bx + c = 0\] To the formula:\[\Large \frac{{ - b \pm \sqrt {{b^2} - 4ac} }}{{2a}}\]

OpenStudy (binary3i):

finding the roots of the equation. it is the point where this graph cuts x axis

OpenStudy (anonymous):

so would it be -1 +the square root of37over 12

OpenStudy (anonymous):

thats the answer i came up with

OpenStudy (anonymous):

\[\Large \begin{array}{l} 6{x^2} - 2x - 1 = 0\\ \frac{{2 \pm \sqrt {{2^2} - 4(6 \cdot - 1)} }}{{2 \cdot 6}}\\ = \frac{{2 \pm \sqrt {28} }}{{12}}\\ = \frac{{1 \pm \sqrt 7 }}{6} \end{array}\] Is what I came up with.